Write in Standard Form. Find C: 2x = -4 + x2^
a) -4
b) 4
c) 2
d) -2
e) None of these

Answer by faceoff57(108)

2x = -4 + x2^
2(-4) = -4 + (-4)^2
-8 = -4 + 16
-8 = 12
2(4) = -4 + (4)^2
8 = -4 + 16
8 = 12
2(2) = -4 + (2)^2
4 = -4 + 4
4 = 0
2(-2)= -4 + (-2)^2
-4 = -4 + 4
-4 = 0
the answer is e) none of these